OpenStudy (anonymous):

find the amount that results from each. $100 invested at 4% compounded quarterly after a period of 2 years

OpenStudy (anonymous):

A = p(1 + r/n)^nt A = 100(1 + .04/2)^4 A = 100 (1.02)^4 $108.24

OpenStudy (anonymous):

I got the answer 108.29 100*(1+(0.04/4)^8